OpenStudy (anonymous):

How do I solve for y if I have x=1/y? Help!!!

OpenStudy (turingtest):

1) multiply both sides by y 2) divide both sides by x

OpenStudy (anonymous):

I was trying to find the inverse of f(x)=1/x. So I got as fas as x=1/y. Did I do something incorrect?

OpenStudy (anonymous):

what is the definition of an inverse function? maybe this will solve your problem

OpenStudy (anonymous):

a function that undoes the other function?

OpenStudy (anonymous):

Its also one-to-one

OpenStudy (anonymous):

in layman's terms that's a good explanation, but mathematically speaking an inverse function is a function such that if you are given a function f(x), with domain D and range R, and this function is one-to-one (meaning all points in D map to exactly one point in R), then there exists an inverse of f(x) called f(x)^-1, where the domain of f(x) is the range of f(x)^-1 and the range of f(x) is the domain of f(x)^-1

OpenStudy (anonymous):

so basically, f(x) 1/x is a one-to-one function... however, if you notice, when you try to solve for it's inverse, it's inverse function has the same domain and range as the regular function. thus, the inverse of y = 1/x is y = 1/x

OpenStudy (anonymous):

If h(x) is the inverse of my f(x) then h'(3) would be solve by?

OpenStudy (anonymous):

just plug in 3 if h(x) = 1/x, and you are solving for h(3), then the answer should be fairly clear

OpenStudy (anonymous):

i was solving for the derivative of h not h. But thanks that does make sense!!